Solve the following question step by step in detail.
∫²₀ xcosx dx
1
Expert's answer
2020-10-14T18:08:20-0400

"\large\int^2_0 xcosx\,dx ="


Using Integration by parts formula,

"\large\int Udv = UV - \int Vdu"


Comparing "\large\int^2_0 xcosx\,dx" and "\large\int Udv"


"U = x" and "dv = cosx \,dx"


Since,

"U = x, \quad \dfrac{du}{dx} = 1"


"\therefore du = dx"


"[U = x\,\,, du =dx]"




Since,

"dv = cosx\,dx, \quad \dfrac{dv}{dx}= cosx"

Integrating,

"\therefore V = sinx"


"[V =sinx\,\,, dv = cosx\,dx]"




"\begin{aligned}



\end{aligned}" "\begin{aligned}

from, \int Udv &= UV - \int Vdu\\\\
\int^a_b Udv &= {\large[UV]^a_b} - \int^a_b Vdu\\\\
\int^2_0xcosx\,dx &= [xsinx]^a_b- \int^2_0sinx\,dx\\\\
\int^2_0xcosx\,dx &= {\large[xsinx]^2_0}- {\large[-cosx]^2_0}\\\\
\int^2_0xcosx\,dx &= [(2sin(2))-(2sin(0))]- [(-cos(2))-(-cos(0))]\\\\
\int^2_0xcosx\,dx &= [2sin2-(0)]- [-cos2-(-1)]\\\\
\int^2_0xcosx\,dx &= 2sin2- [-cos2+1]\\\\
\int^2_0xcosx\,dx &= 2sin2 +cos2 -1

\end{aligned}"



"\Large\therefore \int^2_0xcosx\,dx\space=\space 2sin2 +cos2 -1 \space \approx \space0.4024"